OpenStudy (anonymous):

derivative of (3x)(4x) derivative of (3x)(4x) @Mathematics

OpenStudy (anonymous):

24x

OpenStudy (anonymous):

=f'g + fg'= 12x+12x= 24x

OpenStudy (anonymous):

take the log \[\ln(3x\times 4x)=\ln(3x)+\ln(4x)=\ln(3) + \ln(x)+\ln(4)+\ln(x)\] take the derivative get \[\frac{1}{x}+\frac{1}{x}=\frac{2}{x}\] then multiply by the original function get \[(3x)(4x)\times \frac{2}{x}\] you can simplify if you want

OpenStudy (anonymous):

woah, thats too complicated for me

OpenStudy (anonymous):

satelitte thats wrong, you have to use the product rule

OpenStudy (anonymous):

lagrange trust me

OpenStudy (anonymous):

that is why you need logarithmic differentiation

OpenStudy (anonymous):

but thank you everyone, my Homework is complete

OpenStudy (anonymous):

oh really, i satellite wrong?

OpenStudy (anonymous):

is*

OpenStudy (anonymous):

ya, he didnt realize it is a product rule question

OpenStudy (zarkon):

lol

OpenStudy (anonymous):

make the first time f and the second g, the rule is f ' g + f g ' = derivative of f*g

OpenStudy (anonymous):

term*

OpenStudy (anonymous):

you can also write it as \[e^{\ln(3x\times 4x)}\] and use the chain rule

OpenStudy (anonymous):

oh, satellite how could you forget this was a product rule problem

OpenStudy (anonymous):

ya but he won't know that

OpenStudy (anonymous):

how did you not realize it was too, it is your homework no? so i would assume you just learned it?